# When Does Learning Renormalize? Sufficient Conditions for Power–Law Spectral Dynamics

### v1

@@ 3.2 Statement of sufficient conditions
Condition 4 requires the absence of any intrinsic absolute scale in the renormalized spectral couplings beyond relative separations in logarithmic scale. Unlike Conditions 13, this condition does not follow from generic stability or locality considerations and constitutes a genuinely nontrivial structural requirement.
@@ 3.3 Main theorem
Fix a learning configuration and consider the induced GRSD shell dynamics defined on logarithmic spectral shells. Suppose Conditions 14 hold on a training horizon $t\in[0,T]$. Assume in addition the standard GRSD structural properties: (i) antisymmetry of intra–shell couplings, so that shell–internal transfers cancel in the energy balance; and (ii) renormalizable closure, whereby inter–shell dynamics can be expressed up to a controlled remainder in terms of boundary fluxes between neighboring shells.
@@
is necessarily of power–law form,
$v(\lambda,t)=c(t)\,\lambda^{a},$
for some exponent $a\in\mathbb{R}$ and scalar coefficient $c(t)$.
@@ A.2 Proof of Theorem 1
    Statistical scale covariance (second order). For any input rescaling factor $s>0$,
    $\mathbb{E}_{x}\!\bigl[J(x,t)J(x,t)^{*}\bigr]\;\approx\;s^{2\kappa}\,S_{s}\,\mathbb{E}_{x}\!\bigl[J(sx,t)J(sx,t)^{*}\bigr]\,S_{s}^{-1},$ (13)
    for some $\kappa\in\mathbb{R}$ and invertible maps $S_{s}$ (acting on the output/function space).
@@
Lemma 1 (Time–rescaling covariance of gradient flow).
Let $\theta(t)$ solve the gradient flow
$\dot{\theta}(t)=-\nabla_{\theta}\mathcal{L}(\theta).$
For any $\alpha>0$, define $\mathcal{L}_{\alpha}:=\alpha\,\mathcal{L}$. Then $\theta_{\alpha}(t):=\theta(\alpha t)$ solves the gradient flow associated with $\mathcal{L}_{\alpha}$.
@@
Direct differentiation gives:
$\frac{d}{dt}\theta(\alpha t)=\alpha\dot{\theta}(\alpha t)=-\alpha\nabla\mathcal{L}(\theta(\alpha t))=-\nabla(\alpha\mathcal{L})(\theta_{\alpha}(t)).$ (27)
In the present setting, rescaling the spectral variable $\lambda\mapsto a\lambda$ corresponds, on the intermediate window, to a homogeneous rescaling of the Jacobian–induced quadratic statistics. By Lemma 1, such a rescaling is equivalent to a reparameterization of time $t\mapsto a^{k}t$ for some exponent $k$, depending on the loss and normalization conventions. Consequently, there exist exponents $\gamma_{\varepsilon},\gamma_{J}$ such that
$\varepsilon(a\lambda,t)=a^{\gamma_{\varepsilon}}\,\varepsilon(\lambda,a^{k}t),\qquad J(a\lambda,t)=a^{\gamma_{J}}\,J(\lambda,a^{k}t),$
and hence
$v(a\lambda,t)=a^{\gamma_{J}-\gamma_{\varepsilon}}\,v(\lambda,a^{k}t).$
Equation (A.2) is a functional equation encoding simultaneous log–shift invariance and gradient–flow covariance. Writing $a=e^{\tau}$ and $\lambda=e^{s}$, it follows that
$v(s+\tau,t)=e^{(\gamma_{J}-\gamma_{\varepsilon})\tau}\,v(s,e^{k\tau}t).$
Condition 4 rules out the appearance of a time–dependent reference origin in $s$. Indeed, if $v$ depended nontrivially on the combined variable $s+k^{-1}\log t$, then the renormalized closure would select an absolute log–scale depending on $t$, contradicting log–shift invariance on the intermediate window [wilson1983renormalization, kadanoff1966scaling, forster1977large]. Therefore the tt–dependence in (A.2) must factor out, yielding
$v(\lambda,t)=c(t)\,\lambda^{a},\qquad a:=\gamma_{J}-\gamma_{\varepsilon}.$

### v3

@@ 3.2 Statement of sufficient conditions
Condition 4 requires the absence of any intrinsic absolute scale in the renormalized spectral couplings beyond relative separations in logarithmic scale. It forces that $\dot{s}=\frac{\dot{\lambda}}{\lambda}$ is invariant to the absolute scale of $\lambda$. Unlike Conditions 13, this condition does not follow from generic stability or locality considerations and constitutes a genuinely nontrivial structural requirement.
@@ 3.3 Main theorem
Fix a learning configuration and consider the induced GRSD shell dynamics defined on logarithmic spectral shells. Suppose Conditions 14 hold on a training horizon $t\in[t_{0},T]$. Assume in addition the standard GRSD structural properties: (i) antisymmetry of intra–shell couplings, so that shell–internal transfers cancel in the energy balance; and (ii) renormalizable closure, whereby inter–shell dynamics can be expressed up to a controlled remainder in terms of boundary fluxes between neighboring shells.
@@
then the renormalized GRSD velocity field is uniquely constrained to
$v(\lambda,t)=\frac{c_{0}}{t}\lambda,t\in[t_{0},T]$
for a scalar coefficient $c_{0}=\frac{v(\lambda,t_{0})}{\lambda}$.
Remark 1 (Effective time and learning–rate schedules).
Throughout this work, the time variable $t$ denotes the effective time associated with the continuous–time gradient flow
$\dot{\theta}(t)=-\nabla_{\theta}\mathcal{L}(\theta(t)),$
which corresponds to a constant learning rate and no explicit scheduler. All covariance and scaling properties used in our analysis are derived directly from this ODE.
When a learning–rate schedule $\eta(t)$ is present, the parameter dynamics take the form $\dot{\theta}=-\eta(t)\nabla_{\theta}\mathcal{L}$ and can be mapped to the above gradient–flow form by introducing a reparameterized time lifunctional
$\tau(t)=\int_{0}^{t}\eta(s)\,ds.$
All results in this paper then apply with $t$ replaced by the effective time $\tau$.
@@ A.2 Proof of Theorem 1
    Log–shift invariance of renormalized shell couplings. On an intermediate spectral window, the renormalized shell coupling statistics are translation invariant in the logarithmic coordinate $s=\log\lambda$, i.e. there exists a kernel $K_{h}(\Delta)$, $\Delta=(j-i)h$, such that
    where $\mathrm{err}(n,h,L)\to 0$ in the joint limit of large width $n$, small bin size $h$, and sufficient depth $L$.
@@
Lemma 1 (Gradient-flow covariance fixes the time dependence of $v$).
Consider gradient-flow training with mean-squared error loss
$\mathcal{L}(\theta)=\tfrac{1}{2}\|e(\theta)\|^{2},\qquad\dot{\theta}(t)=-\nabla_{\theta}\mathcal{L}(\theta(t)),$
and induced error dynamics
$\dot{e}(t)=-M(t)e(t),\qquad M(t)=J_{\theta(t)}J_{\theta(t)}^{*}.$
Let $\lambda(t)$ denote an eigenvalue of $M(t)$ and define the spectral drift
$v(\lambda,t):=\frac{d\lambda}{dt}.$
Then $v$ satisfies the covariance functional equation
$v(a\lambda,t/a)=a^{2}\,v(\lambda,t),\qquad\forall a>0,$ (26)
and consequently admits the representation
$v(\lambda,t)=t^{-2}F(\lambda t),\qquad F(u):=v(u,1),\qquad t>0.$ (27)
@@
Step 1: Loss rescaling induces time reparameterization.
For any $a>0$, consider the rescaled loss $\mathcal{L}_{a}:=a\,\mathcal{L}$. The corresponding gradient flow satisfies
$\dot{\theta}_{a}(t)=-\nabla_{\theta}\mathcal{L}_{a}(\theta_{a}(t))=-a\,\nabla_{\theta}\mathcal{L}(\theta_{a}(t)).$
If $\theta(t)$ solves the original gradient flow, then $\theta_{a}(t):=\theta(at)$ satisfies
$\dot{\theta}_{a}(t)=a\,\dot{\theta}(at)=-a\,\nabla_{\theta}\mathcal{L}(\theta(at))=-a\,\nabla_{\theta}\mathcal{L}(\theta_{a}(t)),$
and hence solves the rescaled flow. Thus, rescaling the loss by $a$ is equivalent to the time reparameterization $t\mapsto at$.
Step 2: Loss rescaling multiplies the error operator.
For MSE loss, $\nabla_{\theta}\mathcal{L}(\theta)=-J_{\theta}^{*}e$, so $\dot{\theta}=J_{\theta}^{*}e$. Differentiating $e=y-f_{\theta}$ yields
$\dot{e}=-J_{\theta}\dot{\theta}=-J_{\theta}J_{\theta}^{*}e=-M(t)e.$
Under $\mathcal{L}_{a}=a\mathcal{L}$, the parameter flow becomes $\dot{\theta}_{a}=aJ_{\theta}^{*}e$, and the induced error dynamics are
$\dot{e}_{a}=-J_{\theta}\dot{\theta}_{a}=-a\,J_{\theta}J_{\theta}^{*}e_{a}=-a\,M(t)e_{a}.$
Hence, at the level of the error ODE, loss rescaling induces
$M(t)\mapsto a\,M(t),\quad\text{and therefore}\quad\lambda(t)\mapsto a\,\lambda(t).$
Step 3: Covariance of eigenvalue trajectories.
Combining Steps 1–2, the eigenvalue trajectories obey
$\lambda_{a}(t)=a\,\lambda(at),$
where $\lambda_{a}$ denotes the eigenvalue under $\mathcal{L}_{a}$. Differentiating with respect to $t$ gives
$v_{a}(t)=\dot{\lambda}_{a}(t)=\frac{d}{dt}\big(a\,\lambda(at)\big)=a^{2}\,\dot{\lambda}(at)=a^{2}\,v(at).$
Rewriting this relation in field form yields (26).
Step 4: Solving the functional equation.
Fix $t>0$ and choose $a=t$ in (26). Then
$v(t\lambda,1)=t^{2}\,v(\lambda,t),$
which can be rearranged as
$v(\lambda,t)=t^{-2}v(t\lambda,1).$
Defining $F(u):=v(u,1)$ and substituting $u=\lambda t$ yields (27). ∎
Condition 4 rules out that, on an intermediate spectral window, shell couplings are translation–invariant in $s=\log\lambda$. Let $s=\log\lambda$. Concretely, there exists a kernel $K_{h}$ such that
$\widehat{K}_{ij}=K_{h}\!\big((j-i)h\big)+\mathrm{err}(n,h,L),\qquad s_{i}=ih,$
with
$\mathrm{err}(n,h,L)\to 0$
in the joint limit of large width $n$, small log–bin size $h$, and sufficient depth $L$.
As a consequence, any effective drift velocity in $s$-space cannot depend explicitly on the absolute position $s$. In particular,
$u(s,t):=\dot{s}=\frac{1}{\lambda}\dot{\lambda}=\frac{v(\lambda,t)}{\lambda}=c(t)$
is independent of $\lambda$ within the window.
Equivalently,
$\boxed{\;v(\lambda,t)=\lambda\,c(t)\quad\text{on the intermediate spectral window.}\;}$
Combining this formula with $v(\lambda,t)=t^{-2}v(t\lambda,1).$, we have:
$\boxed{\;v(\lambda,t)=c(t_{0})\frac{\lambda}{t}\ \quad\text{on the intermediate spectral window $t\in[t_{0},T]$.}\;}$